Skip to main contentSkip to solution

A bag contains 17 tickets numbered from 1 to 17. A ticket is drawn, and then another ticket is drawn without replacing the first one. Then the probability that both the tickets show at least one even number, is

Solution

✅ Correct Option: 3

There are 8 even and 9 odd tickets. P(no even) =917×816=934=\frac{9}{17}\times\frac{8}{16}=\frac{9}{34}, so P(at least one even) =1−934=2534=1-\frac{9}{34}=\frac{25}{34}.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question